OpenStudy (anonymous):

please i beg Choose the correct conic section to fit the equation. Circle Ellipse Parabola Hyperbola 1. 1/8 x^2 2. x=1/16y^2 and 3.If the equation of a circle is (x - 2)2 + (y - 6)2 = 4, it passes through point (5, 6). True False

OpenStudy (anonymous):

3. Is False I don't understand 1 and 2

OpenStudy (anonymous):

for 1 and 2 you have to say if they are Circle, Ellipse, Parabola, or Hyperbola i dont know these

OpenStudy (anonymous):

for 1 and 2 you have to say if they are Circle, Ellipse, Parabola, or Hyperbola i dont know these

OpenStudy (anonymous):

Both 1 and 2 are parabolas.

Nnesha (nnesha):

for hyperbola x and y both should have square and there should be negative between \[\large\rm \frac{ x^2 }{a^2 } \color{blue}{-} \frac{ y^2 }{ b^2 }=1\] ellipse almost same but there should be PLUS sign instead negative

Nnesha (nnesha):

parabola only one variable should have x^2 <---exponent

Nnesha (nnesha):

square *

Nnesha (nnesha):

\[\huge\rm x^2 =4p(y-k)\]parabola

OpenStudy (anonymous):

oh i see!

Nnesha (nnesha):

:-)